Determine whether the graphs of the two equations are perpendicular
y=8x-8
2y=5-x

Answer by nyc_function(2741)

If they are penpendicular, then the product of their slopes will be -1.
y = 8x - 8
2y = 5 - x
We need to write 2y = 5 - x in the slope-intercept form y = mx + b.
To do so, we must isolate y.
2y = 5 - x
y = (5 - x)/2
y = 5/2 -(1/2)(x)
The slope is the number next to x.
As you can see, the slopes are not the same nor do they yield -1 when multiplied. So, the two given equations are not perpendicular.